OpenStudy (anonymous):

In the triangle below, what ratio is sin θ? answers: A. 12/13 B. 13/12 C. 13/5 D. 12/5

OpenStudy (anonymous):

I know that sin is opp/hyp

OpenStudy (anonymous):

true what does the triangle look like?

OpenStudy (anonymous):

a right triangle

OpenStudy (anonymous):

lol

OpenStudy (anonymous):

yeah but it mean where in the triangle is \(\theta\) ?

OpenStudy (anonymous):



OpenStudy (anonymous):

my bad

OpenStudy (anonymous):

oh actually it doesn't matter, since the only one that has the hypotenuse in the denominator is \(\frac{12}{13}\)

OpenStudy (anonymous):

ok good, so you have "opposite over hypotenuse" is \(\frac{36}{39}\) reduce

OpenStudy (anonymous):

I am but it doesn't give me the right answer :o

OpenStudy (anonymous):

36/2 = 18 39/2 = 19.5

OpenStudy (anonymous):

try dividing by the common factor of 3 rather than 2

OpenStudy (anonymous):

ohhhh 36/3 = 12 39/3 = 13 so 12/13 ? (:

OpenStudy (anonymous):

yeah, since 39 is not even dividing by 2 will not help your cause

OpenStudy (anonymous):

okay I got it thank you!

OpenStudy (anonymous):

yw